OpenStudy (anonymous):

Solve D = 1/4 (e – 3f) for e. You must show all work.

jimthompson5910 (jim_thompson5910):

D = 1/4 (e – 3f) 4D = e – 3f 4D + 3f = e So e = 3f + 4D

OpenStudy (anonymous):

D=1/4(e-3f) \[D(4) = e-3f\] \[4d+3f=e\]